OpenStudy (anonymous):

What is the solution to the system? Line 1 x y –1 9 3 –3 Line 2 x y –1 –1 2 5 ( , ) @Directrix plz help

Directrix (directrix):

Go here and crank out the equations: http://www.mathportal.org/calculators/analytic-geometry/two-point-form-calculator.php

OpenStudy (anonymous):

y-slope solution right?

Directrix (directrix):

That is okay. I'm waiting for you to post the equation for #1 and then #2

OpenStudy (anonymous):

line 1= y=-3x+6 line 2= y=2x+1

Directrix (directrix):

That is what I got. Now, go here: http://www.wolframalpha.com/ Enter the two equations in this form: y = -3x + 6, y = 2x + 1 and click the = tab.

Directrix (directrix):

You'll see the graphs and the point of intersection which is the solution.

OpenStudy (anonymous):

x=1 y=3

OpenStudy (anonymous):

thats what i got

Directrix (directrix):

That is what I got, too.

OpenStudy (anonymous):

thanks again your a life saver

OpenStudy (anonymous):

can you help with more?

Directrix (directrix):

It is easy to get the points mixed up in that equation of a line calculator so double-check that each time you use it.

OpenStudy (anonymous):

yea

Directrix (directrix):

Yes, close this thread and open a new one.

OpenStudy (anonymous):

ok see ya there